How can I solve the expression x^(p/q)...for example....5^(3/2),7^(1/3)...8^(4/5)....

Grade:11

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear sundeep

x^(p/q) has q solution if you take complex value as well

here i will consider only real value

you can write x^(p/q)  as (xp)1/q

                                             or( x1/q)p

                             for example   (125)2/3

                                                    =(1251/3)2

                                                    =52

                                                    =25

           or you can solve it by using log

         take y=x^(p/q)

               logy=p/q log x

                       = p/q * some value

                       = some value

        take anti log

                   y= some value

 


Please feel free to post as many doubts on our discussion forum as you can. If you find any question Difficult to understand - post it here

and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin